### v1

@@ 6 Conclusions and discussion
In this paper, we introduced two bounded-confidence models (BCMs) of multidimensional opinions: a synchronous model extending the Hegselmann–Krause (HK) model and an asynchronous model extending the Deffuant–Weisbuch (DW) model. We interpret the multidimensional opinion vectors in our models as opinions on multiple related topics and we defined topic-weighted discordance functions to quantify the distance between opinion vectors. These topic-dependent discordance functions depend not only on the opinions for a specific topic $k$ but also incorporate the influence of other topics. For each topic $k$, the discordance function $d_{k}$ calculates a weighted average of the opinion differences across all topics, with a topic weight parameter $\omega$ determining how much more weight is assigned to the opinion difference in topic $k$ than the opinion difference in other topics.
@@
We studied our topic-weighted BCMs by examining the convergence and steady-state properties of the dynamics. We proved that the limit of the opinion vectors in our BCMs exists as time goes to infinity and the limit opinions of two adjacent nodes are either the same or separated in the sense that their topic-weighted discordances are greater than or equal to the confidence bound. To characterize the separation of distinct limit opinions, we analyzed regions of receptiveness to rigorously quantify the neighborhood in the opinion space of a steady-state opinion cluster where no other opinion cluster exists. We also defined time-dependent effective graphs and showed their convergence to the effective graphs associated with the steady-state opinions. We used these effective graphs to establish the stopping criterion for our numerical simulations.
@@
In addition to investigating the theoretical aspects of our topic-weighted BCMs, we numerically simulated our topic-weighted BCMs with initial opinions sampled from various distributions, including both independent and interdependent initial opinions distributions on different topics. We demonstrated that the choice of initial opinion distribution significantly influences the degree of opinion fragmentation (quantified by the order parameter $Q$) and the BCM parameters — specifically, the topic weight $\omega$ and confidence bound $c$ — at which the transition between consensus and fragmentation occurs. For both our topic-weighted BCMs, for uniform and truncated Gaussian initial opinion distributions with a fixed $c$, decreasing $\omega$ tends to reduce opinion fragmentation, likely due to the expansion of the regions of receptiveness as $\omega$ decreases. However, for the wedge distribution, varying $\omega$ for fixed $c$ has little impact on opinion fragmentation, indicating that the effect of $\omega$ depends on the initial opinion distribution. In the case of our topic-weighted HK model applied to two-community SBM graphs, the graph structure shows minimal influence on opinion fragmentation when initial opinions are independent of the community. However, fragmentation increases when initial opinions are correlated with community membership. Our simulations highlight the substantial impact of initial opinion distributions on the behavior of our topic-weighted BCMs and baseline BCMs. Given that most research on BCMs only considers uniform initial opinions, we encourage researchers studying BCMs and other models of opinion dynamics with continuous opinions to consider a wider range of initial opinion distributions.
@@ A.1 A complete graph with truncated Gaussian initial opinions
In Fig. 13, we show some plots of the opinion trajectories for simulations with Gaussian $G(0.22,0)$ initial opinions and confidence bound $c=0.075$ for both our topic-weighted HK model with topic weight $\omega=0.1$ and the baseline HK model. As was the case for uniform initial opinions (see Fig. 6), for the Gaussian initial opinions, simulations of the baseline HK model (see Fig. 13B) tend to result in the initial opinions of the nodes in each final opinion cluster appearing to form a rectangular region containing the corresponding mean final opinion. Simulations of our topic-weighted HK model with $\omega=0.1$ (see Fig. 13A) tend to result in the initial opinions of nodes in each final opinion cluster forming non-rectangular shapes. For the Gaussian $G(0.22,0)$ initial opinions, nodes that have initial opinions that are away from $(0.5,0.5)$, tend to have opinions that move towards $(0.5,0.5)$ over time. In Fig. 13, for final opinion clusters with corresponding mean final opinion not near $(0.5,0.5)$, the mean final opinion is near the boundaries of the region formed by the corresponding initial opinions. Intuitively, this makes sense because the Gaussian $G(0.22,0)$ distribution has a larger probability of initial opinions near the distribution mean at $(0.5,0.5)$. If there are more nodes with initial opinions near $(0.5,0.5)$, these nodes will have more weight in HK update rule (see (7)) and will pull the opinions of other nodes towards the center of the opinion space at $(0.5,0.5)$.
@@
In Fig. 14, we show some plots of the opinion trajectories for simulations with Gaussian $G(0.22,0.8)$ initial opinions and $c=0.075$ for both our topic-weighted HK model with $\omega=0.1$ and the baseline HK model. As was the case for uniform and $G(0.22,0)$ initial opinions, the final opinion clusters in simulations of the baseline HK model tend to have corresponding initial opinions that form rectangular regions. The final opinion clusters in simulations of our topic-weighted HK model with $\omega=0.1$ tend to have corresponding initial opinions that form non-rectangular regions. In the simulation of our topic-weighted HK model with $\omega=0.1$ (see Fig. 13A), there are two nodes that are each in a final opinions cluster consisting of just themselves; they have final opinions near $(0.33,0.52)$ and $(0.91,0.37)$. These two nodes have initial opinions away from the skew axis (i.e., the line $x^{1}=x^{2}$) of the $G(\sigma=0.22,\rho=0.8)$ distribution and quickly are unreceptive to all other nodes. There are four other final opinion clusters that have the 1,998 other nodes in the graph. These four final opinion clusters have mean final opinions near the skew axis of the $G(0.22,0.8)$ distribution. Which final opinion cluster a node belongs to appears to relate to which point on the skew axis their initial opinion is closest to.
@@ A.2 SBM graphs with Gaussian initial opinions
In Fig. 15, we show the final order parameters $Q(T_{f})$ (see (70)) of our simulations of our topic-weighted HK model on two-community SBM graphs with truncated Gaussian $G(0.22,0)$ initial opinions for various values of BCM parameters (namely, the confidence bound $c$ and topic weight $\omega$). We observe the same trends as our two-community SBM simulations with uniform initial opinions (see section 5.1.2). In particular, for both the community-dependent and randomly-assigned initial opinions, we tend to observe larger $Q(T_{f})$ (and correspondingly less opinion fragmentation) as we either (1) increase $c$ for fixed $\omega$ or; (2) decrease $\omega$ for fixed $c$. Additionally, for fixed $\omega$, the minimum value of $c$ for our simulations to always reach consensus is larger for the community-dependent initial opinions (see Fig. 15A) compared to the randomly-assigned initial opinions Fig. 15B).
@@ Appendix B Additional results of our topic-weighted DW simulations
In Fig. 17, we show some plots of the opinion trajectories for simulations of our topic-weighted DW model with $\omega=0.4$ and $c=0.25$ with wedge initial opinions. Each panel shows a simulation with a different set of initial opinions, and we show examples of simulations that have 1–3 final opinion clusters. In our topic-weighted DW model, there is randomness in which pair of nodes and which topic is selected at each time for the update rule (8). Consequently, nodes in our topic-weighted DW model can have opinion trajectories that traverse more of the opinion space than nodes in our topic-weighted HK model (see Fig. 7). Our simulations of our topic-weighted DW model with wedge initial opinions typically have 2–5 final opinion clusters. Randomness in the selection of nodes and topics contribute to how many final opinion clusters are in a simulation. As we discussed in section 5.2, for $c=0.25$, there appears to be noticeably less opinion fragmentation for $0.4\leq\omega\leq 0.6$ than for other values of $\omega$. It may be that for $c=0.25$, for the wedge initial opinion distribution, the shape and size of the regions of receptiveness when $0.4\leq\omega\leq 0.6$ make it more likely to have fewer final opinion clusters and smaller final order parameter than other values of $\omega$.
@@ Acknowledgments
We thank Mason A. Porter and Dominic Yang for helpful discussions. We also thank Deanna Needell and Jacob Foster for helpful comments.

### v2

@@ 6 Conclusions and discussion
In this paper, we introduced two bounded-confidence models (BCMs) of multidimensional opinions: a synchronous model extending the Hegselmann–Krause (HK) model and an asynchronous model extending the Deffuant–Weisbuch (DW) model. We interpret the multidimensional opinion vectors in our models as opinions on multiple related topics, and we defined topic-weighted discordance functions to quantify the distance between opinion vectors. These topic-dependent discordance functions depend not only on the opinions for a specific topic $k$ but also incorporate the influence of other topics. For each topic $k$, the discordance function $d_{k}$ calculates a weighted average of the opinion differences across all topics, with a topic weight parameter $\omega$ determining how much more weight is assigned to the opinion difference in topic $k$ than the opinion difference in other topics.
@@
We studied our topic-weighted BCMs by examining the convergence and steady-state properties of the dynamics. We proved that the limit of the opinion vectors in our BCMs exists as time goes to infinity, and the limit opinions of two adjacent nodes are either the same or separated in the sense that their topic-weighted discordances are greater than or equal to the confidence bound. To characterize the separation of distinct limit opinions, we analyzed regions of receptiveness to rigorously quantify the neighborhood in the opinion space of a steady-state opinion cluster where no other opinion cluster exists. We also defined time-dependent effective graphs and showed their convergence to the effective graphs associated with the steady-state opinions. We used these effective graphs to establish the stopping criterion for our numerical simulations.
@@
In addition to investigating the theoretical aspects of our topic-weighted BCMs, we numerically simulated our topic-weighted BCMs with initial opinions sampled from various distributions, including both independent and interdependent initial opinions distributions on different topics. We demonstrated that the choice of initial opinion distribution significantly influences the degree of opinion fragmentation (quantified by the order parameter $Q$) and the BCM parameters — specifically, the confidence bound $c$ and topic weight $\omega$ — at which the transition between consensus and fragmentation occurs. For both our topic-weighted BCMs, for uniform and truncated Gaussian initial opinion distributions with a fixed $c$, decreasing $\omega$ tends to reduce opinion fragmentation, likely due to the expansion of the regions of receptiveness as $\omega$ decreases. However, for the wedge distribution, varying $\omega$ for fixed $c$ has little impact on opinion fragmentation, indicating that the effect of $\omega$ depends on the initial opinion distribution. In the case of our topic-weighted HK model applied to two-community SBM graphs, the graph structure shows minimal influence on opinion fragmentation when initial opinions are independent of the community. However, fragmentation increases when initial opinions are correlated with community membership. Our simulations highlight the substantial impact of initial opinion distributions on the behavior of our topic-weighted BCMs and baseline BCMs. Given that most research on BCMs only considers uniform initial opinions, we encourage researchers studying BCMs and other models of opinion dynamics with continuous-valued opinions to consider a wider range of initial opinion distributions.
In our numerical simulations of our topic-weighted BCMs, we observed that the shape of the region of receptiveness, which is determined by $\omega$ in our models, can affect the amount of opinion fragmentation and the qualitative behavior of the opinion trajectories. Most prior studies on BCMs with multidimensional opinions consider Euclidean distance [40, 38, 23, 9, 21, 30, 12]. The Euclidean norm yields spherical regions of receptiveness, which differ from the polytope-shaped regions of receptiveness in our models. Multidimensional opinion models have a variety of choices for modeling opinion distance, which yield different regions of receptiveness. As an area of future work, it is of interest to compare the opinion dynamics of BCMs with different opinion distances and regions of receptiveness.
@@ A.1 A complete graph with truncated Gaussian initial opinions
In Fig. 13, we show some plots of the opinion trajectories for simulations with Gaussian $G(0.22,0)$ initial opinions and confidence bound $c=0.075$ for both our topic-weighted HK model with topic weight $\omega=0.1$ and the baseline HK model. As was the case for uniform initial opinions (see Fig. 6), for the Gaussian initial opinions, simulations of the baseline HK model (see Fig. 13B) tend to result in the initial opinions of the nodes in each final opinion cluster appearing to form a rectangular region containing the corresponding mean final opinion. Simulations of our topic-weighted HK model with $\omega=0.1$ (see Fig. 13A) tend to result in the initial opinions of nodes in each final opinion cluster forming non-rectangular shapes. We observe that nodes with initial opinions that are away from the distribution mean at $(0.5,0.5)$ tend to have opinions that move towards $(0.5,0.5)$ over time. In Fig. 13, for final opinion clusters with corresponding mean final opinion not near $(0.5,0.5)$, the mean final opinion is near the boundary of the region formed by the corresponding initial opinions. Intuitively, this makes sense because the Gaussian $G(0.22,0)$ distribution has a larger probability of initial opinions near the distribution mean at $(0.5,0.5)$. If there are more nodes with initial opinions near the distribution mean, then these nodes have more weight in the HK update rule (see (7)) and pull the opinions of other nodes towards the distribution mean.
@@
In Fig. 14, we show some plots of the opinion trajectories for simulations with Gaussian $G(0.22,0.8)$ initial opinions and $c=0.075$ for both our topic-weighted HK model with $\omega=0.1$ and the baseline HK model. As was the case for uniform and $G(0.22,0)$ initial opinions, we tend to observe rectangular and non-rectangular regions formed by the final opinion clusters in the opinion-trajectory plots for the baseline and topic-weighted HK models, respectively. In the simulation of our topic-weighted HK model with $\omega=0.1$ (see Fig. 13A), there are two nodes that are each in a final opinions cluster consisting of just themselves; they have final opinions near $(0.33,0.52)$ and $(0.91,0.37)$. These two nodes have initial opinions away from the skew axis (i.e., the line $x^{1}=x^{2}$) of the $G(0.22,0.8)$ distribution and quickly become unreceptive to all other nodes. There are four other final opinion clusters that have the 1,998 other nodes in the graph. These four final opinion clusters have mean final opinions near the skew axis of the $G(0.22,0.8)$ distribution. Which final opinion cluster a node belongs to appears to relate to which point on the skew axis their initial opinion is closest to.
@@ A.2 SBM graphs with Gaussian initial opinions
In Fig. 15, we show the final order parameters $Q(T_{f})$ (see (70)) of our simulations of our topic-weighted HK model on two-community SBM graphs with truncated Gaussian $G(0.22,0)$ initial opinions for various values of BCM parameters (namely, the confidence bound $c$ and topic weight $\omega$). We observe the same trends as our two-community SBM simulations with uniform initial opinions (see section 5.1.2). In particular, for both the community-dependent and randomly-assigned initial opinions, we tend to observe larger $Q(T_{f})$ (and correspondingly less opinion fragmentation) as we either (1) increase $c$ for fixed $\omega$; or (2) decrease $\omega$ for fixed $c$. Additionally, for fixed $\omega$, the minimum value of $c$ for our simulations to always reach consensus is larger for the community-dependent initial opinions (see Fig. 15A) compared to the randomly-assigned initial opinions Fig. 15B).
@@ Appendix B Additional results of our topic-weighted DW simulations
In Fig. 17, we show some plots of the opinion trajectories for simulations of our topic-weighted DW model with $\omega=0.4$ and $c=0.25$ with wedge initial opinions. Each panel shows a simulation with a different set of initial opinions, and we show examples of simulations that have 1–3 final opinion clusters. In our topic-weighted DW model, there is randomness in which pair of nodes and which topic is selected at each time for the update rule (8). Consequently, nodes in our topic-weighted DW model can have opinion trajectories that traverse more of the opinion space than nodes in our topic-weighted HK model (see Fig. 7). Our simulations of our topic-weighted DW model with wedge initial opinions typically have 2–5 final opinion clusters. Randomness in the selection of nodes and topics contributes to how many final opinion clusters are in a simulation. As we discussed in section 5.2, for $c=0.25$, there appears to be noticeably less opinion fragmentation for $0.4\leq\omega\leq 0.6$ than for other values of $\omega$. It may be that for $c=0.25$, for the wedge initial opinion distribution, the shape and size of the regions of receptiveness when $0.4\leq\omega\leq 0.6$ make it more likely to have fewer final opinion clusters and smaller final order parameter than other values of $\omega$.
@@ Acknowledgments
We thank Mason A. Porter and Dominic Yang for helpful discussions. We also thank Deanna Needell and Jacob Foster for helpful comments. Additionally, we thank our anonymous referee for their helpful comments.